# Learning a Neural Solver for Parametric PDE to Enhance Physics-Informed Methods

### v2

@@ Acknowledgments
We acknowledge the financial support provided by DL4CLIM (ANR-19-CHIA-0018-01), DEEPNUM (ANR-21-CE23-0017-02), PHLUSIM (ANR-23-CE23-0025-02), and PEPR Sharp (ANR-23-PEIA-0008, ANR, FRANCE 2030).
@@ Appendix A Detailed related Work
Many tools for numerically solving PDEs have been developed for years. The standard methods for PDE include Finite Differences (FD), Finite Volume (FV), Finite Element Method (FEM), spectral and multigrid methods, and many others (S. H, 2012; Liu, 2009). While these methods are widely used, they often suffer from a high computational cost for complex problems or high-precision simulations. To address these challenges, integrating deep learning (DL) into solvers has emerged as a promising approach. Current solutions include incorporating correction terms into mathematical solvers to reduce numerical errors (Um et al., 2021). Some work such as Hsieh et al. (2019); Li et al. (2023a); Rudikov et al. (2024); Kopaničáková et al. (2023) build a method to directly enhance the convergence of numerical solvers through preconditioner learning. As an example, Rudikov et al. (2024); Li et al. (2023a) uses a neural operator to approximate conditioner for the flexible conjugate gradient method or Hsieh et al. (2019) for the Jacobi method. Another example of preconditioner learning lies in Li et al. (2023a), where the author uses GNN to assess symmetry and positive definiteness.
@@
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) have been a pioneering work in the development of DL method for physics. In these models, the solution is a neural network that is optimized using the residual loss of the PDE being solved. However, this method suffers from several drawbacks. First, as formulated in Raissi et al. (2019), PINNs can solve one instance of an equation at a time. Any small change in the parameters of the PDE involves a full retraining of the network. Efforts such as Beltrán-Pulido et al. (2022); Zhang et al. (2023) have attempted to address this limitation by introducing parametric versions of PINNs capable of handling parametric equations, while Huang et al. (2022); Qin et al. (2022) explores meta-learning approaches. Other approaches to improve PINNs generalization include using neural operator (Wang et al., 2021) or hyper-network (de Avila Belbute-Peres et al., 2021). Moreover, PINNs have shown convergence difficulties: Krishnapriyan et al. (2021) show that PINNs’ losses have complex optimization landscapes, complicating training despite adequate neural network expressiveness. Approaches like those detailed in Wang et al. (2022) adopt a Neural Tangent Kernel (NTK) perspective to identify reasons for failure and suggest using adaptive weights during training to enhance performance. Additionally, studies such as Ryck et al. (2023) demonstrate that PINNs suffer from ill-conditioned losses, resulting in slow convergence of gradient descent algorithms. Recently, (Rathore et al., 2024) show how training strategies can improve convergence of Physics-Informed Neural Networks and show that specific optimizers such as L-BFGS acts as conditioner on the physical losses.
@@
In contrast to the unsupervised training of Physics-informed Neural Networks, purely data-driven models have demonstrated remarkable capabilities for PDE simulation and forecasting. In most of the existing literature, the entire solver is replaced by a DL architecture and focuses on directly computing the solution from a given input data. A widely studied setting is operator learning which learns mappings between function spaces (Li et al., 2020; Kovachki et al., 2023; Lu et al., 2021). This method is very efficient, with the downside of relying on quite large quantities of data for training in order to ensure adequate generalization. Additionally, the neural network does not have access to the PDE in itself, only indirectly through the data. To ensure physical constraint in purely-data-driven training, hybrid models have been proposed. The latter rely on both the available physical knowledge and some available data. Some examples include the Aphinity model (Yin et al., 2021) (where the authors have partial knowledge of the physics and learned the remaining dynamics from data), Physics-informed Deep Operator Networks (PIDON) (Wang et al., 2021; Goswami et al., 2022), Physics-informed Neural Operator (PINO) (Li et al., 2023b) (DeepONet architecture (Lu et al., 2021) or Neural Operator models (Kovachki et al., 2023; Li et al., 2020) respectively with a combination of data and physical losses).
Learning to solve
@@
: Improving the learning scheme and optimizers through data-driven training has been studied since Li & Malik (2016) and Andrychowicz et al. (2016). These works propose to learn the optimizer of neural networks, which are classically optimized through gradient-based algorithms such as Adam. They focus on improving training strategies for neural networks, which do not suffer from the optimization issues and ill-conditioning properties of physics-informed losses. We refer the reader to the survey of Chen et al. (2021) for a complete overview. The closer work to ours is the very recent work of Bihlo (2024) in which the author asseses the capabilities of learned optimizers for physics-informed neural networks. The main difference with our work relies on the problem setting. Bihlo (2024) considers learning an optimizer on a single equation, and for different neural networks initialization, while we focus on efficiently solving several instances of parametric PDE with varying PDE parameters $\gamma,f,g$.
Appendix B Proof of convergence
In this section, we propose a clarification on the convergence properties of Physics-Informed Neural Networks (PINNs) when a neural network is introduced into the framework. In section 2, we showed that gradient descent on the PINNs loss for a linear equation guarantees convergence, with the convergence rate governed by the condition number $\kappa(A)$. Here, we will address how incorporating a neural network impacts convergence, particularly focusing on the condition number of the system. We use the notations and assumptions introduced in section 2.
B.1 Convergence in the Linear Case
In the linear case, the convergence of gradient descent is characterized by the number of steps $N(\epsilon)$ required to achieve an error less than $\epsilon$. This is linearly dependent on $\kappa(A)$ (section 2):
$N(\epsilon)=O\left(\kappa(A)\log\left(\frac{1}{\epsilon}\right)\right)$
@@ B.1 Convergence in the Linear Case
- $A$: The matrix arising from the problem formulation in the basis.
- $\kappa(A)$: The condition number of $A$, which can be quite large, indicating a potential need for many iterations.
B.2 Neural Network Approximation
When introducing a neural network $\mathcal{F}_{\varrho}$, we recognize the challenges in proving convergence due to the complex nature of neural networks. However, a useful approach is to consider the neural network’s behavior through its linear approximation, which allows us to use similar analysis techniques.
Linearization Approach
The neural network $\mathcal{F}_{\varrho}$ can be approximated by its tangent linear model, $\mathcal{F}_{\varrho}(v)\approx Pv$, where $P=\text{Jac}_{0}(\mathcal{F}_{\varrho})$, its Jacobian. The gradient descent update rule in this case becomes:
$\Theta_{l+1}=\Theta_{l}-\eta P\nabla_{\Theta}\mathcal{L}_{\text{PDE}}(\Theta_{l})=(\text{Id}-\eta PA)\Theta_{l}+c$
Here, $c$ is a constant vector independent of $\Theta$.
Convergence Analysis with Linearization
By considering the linearized model, the updates can be expressed as:
$\Theta_{l+1}=(\text{Id}-\eta PA)^{l+1}(\Theta_{0}-\Theta^{*})+\Theta^{*}$
where:
@@ B.2 Neural Network Approximation
- $\Theta^{*}=\Theta_{0}+A^{-1}c$ is the optimal parameter vector.
This expression parallels the convergence analysis of the original linear system, with the key difference being the introduction of the matrix product $PA$.
Condition Number of $PA$
In this formulation, convergence is governed by the condition number $\kappa(PA)$, rather than $\kappa(A)$:
$N^{\prime}(\epsilon)=O\left(\kappa(PA)\log\left(\frac{1}{\epsilon}\right)\right)$
The introduction of $P$ as a preconditioner often results in $\kappa(PA)<\kappa(A)$. This implies that:
  • the neural network can effectively reduces the condition number, enhancing convergence speed and stability.
We show below that the condition number is improved under some optimality conditions.
B.3 Optimality and Convergence
The neural network $\mathcal{F}_{\varrho}$ is trained to minimize the $L_{2}$ norm $\|u_{L}-u^{*}\|_{2}$, where $u_{L}$ is the output after $L$ iterations and $u^{*}$ is the true solution. At convergence, $u_{L}=u^{*}$, implying $\Theta_{L}=\Theta^{*}$. We can consider the concatenated matrices $\Xi_{0}-\Xi^{*}$ where each row is associated to a sample. Incorporating this optimality condition, assuming the training samples span the parameter space completely (i.e. the matrix $(\Xi-\Xi^{*})^{\top}(\Xi-\Xi^{*})$ is full rank), we have that:
$(\text{Id}-\eta PA)^{l+1}(\Xi_{0}-\Xi^{*})=0$
With $\Xi_{0}\neq\Xi^{*}$, this condition implies that $(\text{Id}-\eta PA)^{l+1}=0$, indicating that $(\text{Id}-\eta PA)$ is nilpotent. As a consequence, all eigenvalues of $\eta PA$ are 1, leading to $\kappa(PA)=1$.
Under these conditions:

### v3

@@ Acknowledgments
We acknowledge the financial support provided by DL4CLIM (ANR-19-CHIA-0018-01), DEEPNUM (ANR-21-CE23-0017-02), PHLUSIM (ANR-23-CE23-0025-02), and PEPR Sharp (ANR-23-PEIA-0008”, “ANR”, “FRANCE 2030”).
@@ Appendix A Detailed related Work
Many tools for numerically solving PDEs have been developed for years. The standard methods for PDE include Finite Differences (FD), Finite Volume (FV), Finite Element Method (FEM), spectral and multigrid methods, and many others (S. H, 2012; Liu, 2009). While these methods are widely used, they often suffer from a high computational cost for complex problems or high-precision simulations. To address these challenges, integrating deep learning (DL) into solvers has emerged as a promising approach. Current solutions include incorporating correction terms into mathematical solvers to reduce numerical errors (Um et al., 2021). Some work such as Hsieh et al. (2019); Li et al. (2023a); Rudikov et al. (2024); Kopaničáková et al. (2023) build a method to directly enhance the convergence of numerical solvers through preconditioner learning. As an example, Rudikov et al. (2024); Li et al. (2023a) use a neural operator to approximate conditioner for the flexible conjugate gradient method or Hsieh et al. (2019) for the Jacobi method. Another example of preconditioner learning lies in Li et al. (2023a), where the author uses GNN to assess symmetry and positive definiteness.
@@
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) have been a pioneering work in the development of DL method for physics. In these models, the solution is a neural network that is optimized using the residual loss of the PDE being solved. However, this method suffers from several drawbacks. First, as formulated in Raissi et al. (2019), PINNs can solve one instance of an equation at a time. Any small change in the parameters of the PDE involves a full retraining of the network. Efforts such as Beltrán-Pulido et al. (2022); Zhang et al. (2023); Cho et al. (2024) have attempted to address this limitation by introducing parametric versions of PINNs capable of handling parametric equations, while Huang et al. (2022); Qin et al. (2022) explores meta-learning approaches. Other approaches to improve PINNs generalization include using neural operators (Wang et al., 2021b), or hyper-network (de Avila Belbute-Peres et al., 2021). Moreover, PINNs have shown convergence difficulties: Krishnapriyan et al. (2021) show that PINNs’ losses have complex optimization landscapes, complicating training despite adequate neural network expressiveness. Approaches like those detailed in Wang et al. (2022) adopt a Neural Tangent Kernel (NTK) perspective to identify reasons for failure and suggest using adaptive weights during training to enhance performance. Additionally, studies such as Ryck et al. (2023) demonstrate that PINNs suffer from ill-conditioned losses, resulting in slow convergence of gradient descent algorithms. Recently, (Rathore et al., 2024) shows how training strategies can improve the convergence of Physics-Informed Neural Networks and show that specific optimizers such as L-BFGS act as conditioners on the physical losses.
@@
In contrast to the unsupervised training of Physics-informed Neural Networks, purely data-driven models have demonstrated remarkable capabilities for PDE simulation and forecasting. In most of the existing literature, the entire solver is replaced by a DL architecture and focuses on directly computing the solution from a given input data. A widely studied setting is operator learning which learns mappings between function spaces (Li et al., 2020; Kovachki et al., 2023; Lu et al., 2021). This method is very efficient, with the downside of relying on quite large quantities of data for training in order to ensure adequate generalization. Additionally, the neural network does not have access to the PDE in itself, only indirectly through the data. To ensure physical constraints in purely data-driven training, hybrid models have been proposed. The latter relies on both the available physical knowledge and some data. Some examples include the Aphinity model (Yin et al., 2021) (where the authors assume partial knowledge of the physics and learned the remaining dynamics from data), Physics-informed Deep Operator Networks (PIDON) (Wang et al., 2021b; Goswami et al., 2022), Physics-informed Neural Operator (PINO) (Li et al., 2023b) (DeepONet architecture (Lu et al., 2021) or Neural Operator models (Kovachki et al., 2023; Li et al., 2020) respectively with a combination of data and physical losses).
Learning to solve:
@@
Improving the learning scheme and optimizers through data-driven training has been studied since Li & Malik (2016) and Andrychowicz et al. (2016). These works propose to learn the optimizer of neural networks, which are classically optimized through gradient-based algorithms such as Adam. They focus on improving training strategies for neural networks, which do not suffer from the optimization issues and ill-conditioning properties of physics-informed losses. We refer the reader to the survey of Chen et al. (2021) for a complete overview. The closer work to ours is the very recent work of Bihlo (2024) in which the author assesses the capabilities of learned optimizers for physics-informed neural networks. The main difference with our work relies on the problem setting. Bihlo (2024) considers learning an optimizer on a single equation, and for different neural networks initialization, while we focus on efficiently solving several instances of parametric PDE with varying PDE parameters $\gamma,f,g$.
Appendix B Theoretical Analysis of our Method and PINNs
Setting.
We consider the following linear PDE:
$\mathcal{D}u(x)$ $=f(x),\quad x\in\Omega,$ (12)
$u(x)$ $=g(x),\quad x\in\partial\Omega,$
where $\Omega\subset\mathbb{R}^{d}$ is an open bounded domain, $\mathcal{D}$ is a linear differential operator, $f(x)$ is a given function in $\Omega$, and $g(x)$ is a given function on the boundary $\partial\Omega$.
B.1 Theoretical Analysis of PINNs
Our aim is to find an approximate solution $u_{\Theta}(x)$, parameterized by $\Theta\in\mathbb{R}^{N}$, $\Theta=\{\theta_{i}\}_{i=1}^{N}$ that minimizes the loss function:
$\mathcal{L}_{\textnormal{PDE}}(\Theta)=\mathcal{L}_{\textnormal{Res}}(\Theta)+\lambda\mathcal{L}_{\textnormal{BC}}(\Theta),$ (13)
@@ B.1 Theoretical Analysis of PINNs
[…]
and $\lambda>0$ is a regularization parameter balancing the PDE residual and boundary conditions.
We perform gradient descent updates with step size $\eta$. At step $k>0$, updates write as:
$\Theta_{k+1}=\Theta_{k}-\eta\nabla_{\Theta}\mathcal{L}_{\textnormal{PDE}}(\Theta_{k}).$
We establish the following theorem regarding the convergence rate of gradient descent.
Theorem 2 (Convergence rate of PINNs).
Given a linear ansatz $u_{\Theta}(x)=\sum_{i=1}^{N}\theta_{i}\phi_{i}(x)$, the number of steps $N(\varepsilon)$ required to achieve an error $\|\Theta_{k}-\Theta^{\ast}\|_{2}\leq\varepsilon$ satisfies:
$N(\varepsilon)=O\left(\kappa(A)\ln\left(\tfrac{1}{\varepsilon}\right)\right),$ (14)
where $\kappa(A)$ is the condition number of the matrix $A\in\mathbb{R}^{n\times n}$ defined by:
[…]
Proof.
Since $u_{\Theta}(x)=\sum_{i=1}^{N}\theta_{i}\phi_{i}(x)$, we have:
$\frac{\partial u_{\Theta}(x)}{\partial\theta_{i}}=\phi_{i}(x),\quad\frac{\partial(\mathcal{D}u_{\Theta}(x))}{\partial\theta_{i}}=\mathcal{D}\phi_{i}(x).$
The gradient of the residual loss is:
[…]
where $\mathcal{D}\phi(x)$ is the vector with components $\mathcal{D}\phi_{i}(x)$. Similarly, the gradient of the boundary loss is:
[…]
where $\phi(x)$ is the vector of basis functions evaluated at $x$. Therefore, the total gradient is:
$\nabla_{\Theta}\mathcal{L}_{\textnormal{PDE}}(\Theta)=A\Theta-b,$
where the (positive, semi-definite) matrix $A$ and vector $b$ are defined as:
[…]
Thus, the gradient descent update becomes:
$\Theta_{k+1}=\Theta_{k}-\eta(A\Theta_{k}-b).$
@@
Subtracting $\Theta^{\ast}$ (the optimal parameter vector satisfying $A\Theta^{\ast}=b$) from both sides:
$\Theta_{k+1}-\Theta^{\ast}=\Theta_{k}-\Theta^{\ast}-\eta A(\Theta_{k}-\Theta^{\ast}).$
Simplifying:
$\Theta_{k+1}-\Theta^{\ast}=(\text{Id}-\eta A)(\Theta_{k}-\Theta^{\ast}).$
By recursively applying the update rule, we obtain:
$\Theta_{k}-\Theta^{\ast}=(\text{Id}-\eta A)^{k}(\Theta_{0}-\Theta^{\ast}).$
Since $A$ is symmetric positive definite, it has eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$ with $\lambda_{i}>0$. To ensure convergence, we require $0<\eta<\frac{2}{\lambda_{\max}(A)}$. Choosing $\eta=\frac{c}{\lambda_{\max}(A)}$ with $20<c<$ 2, we have:
$1-\eta\lambda_{i}=1-c\frac{\lambda_{i}}{\lambda_{\max}(A)}.$
The spectral radius $\rho$ of $\text{Id}-\eta A$ is:
$\rho=\max\left\{\left|1-c\frac{\lambda_{\min}(A)}{\lambda_{\max}(A)}\right|,\left|1-c\right|\right\}=\max\left\{1-\frac{c}{\kappa(A)},|1-c|\right\},$
where $\kappa(A)=\frac{\lambda_{\max}(A)}{\lambda_{\min}(A)}$ is the condition number of $A$. By choosing $10<c<$ 1, we ensure $|1-c|<1$, and since $\kappa(A)\geq 1$, we have $11-\frac{c}{\kappa(A)}<$ 1. Thus, the convergence factor is:
$\rho=1-\frac{c}{\kappa(A)}.$
Therefore:
$\|\Theta_{k}-\Theta^{\ast}\|_{2}\leq\left(1-\frac{c}{\kappa(A)}\right)^{k}\|\Theta_{0}-\Theta^{\ast}\|_{2}.$
To achieve $\|\Theta_{k}-\Theta^{\ast}\|_{2}\leq\varepsilon$, the number of iterations $N(\varepsilon)$ satisfies:
$N(\varepsilon)\geq\frac{\ln\left(\varepsilon/\|\Theta_{0}-\Theta^{\ast}\|_{2}\right)}{\ln\left(1-\frac{c}{\kappa(A)}\right)}.$
Using the inequality $\ln(1-x)\leq-x$ for $10<x<$ 1, we get:
$N(\varepsilon)\leq\frac{\kappa(A)}{c}\ln\left(\frac{\|\Theta_{0}-\Theta^{\ast}\|_{2}}{\varepsilon}\right).$
Thus:
$N(\varepsilon)=O\left(\kappa(A)\ln\left(\tfrac{1}{\varepsilon}\right)\right).$
We have shown that for a linear ansatz $u_{\Theta}(x)=\sum_{i=1}^{N}\theta_{i}\phi_{i}(x)$, the convergence rate of gradient descent depends linearly on the condition number $\kappa(A)$ of the system matrix $A$. A large condition number impedes convergence, requiring more iterations to achieve a desired accuracy $\varepsilon$.
B.2 Theoretical Analysis of our Method
In practice, we often work with multiple data points. For each data point, there is an associated parameter vector $\Theta\in\mathbb{R}^{N}$. We are interested in the iterative update where the gradient is transformed by a neural network $\mathcal{F}$:
$\Theta_{l+1}=\Theta_{l}-\eta\mathcal{F}\left(\nabla_{\Theta}\mathcal{L}_{\textnormal{PDE}}(\Theta_{l})\right),$ (17)
where $\Theta_{l}$ represents the parameter vector at iteration $l$. Recall that $\mathcal{F}$ is trained to minimize the loss after $L$ iteration steps for $M$ data points:
$\mathcal{L}_{\text{DATA}}=\frac{1}{m}\sum_{k=1}^{M}\left\|u_{\Theta_{L}^{(k)}}-u^{\ast}_{k}\right\|_{2}^{2},$ (18)
Theorem 3.
(Convergence rate of our method). Given a linear ansatz $u_{\Theta}(x)=\sum_{i=1}^{N}\theta_{i}\phi_{i}(x)$, assume $\mathcal{F}$ behaves like its linearization $P=\left.\tfrac{\partial\mathcal{F}}{\partial v}\right|_{v=0}$. The number of steps $N^{\prime}(\varepsilon)$ required to achieve an error $\|\Theta_{l}-\Theta^{\ast}\|_{2}\leq\varepsilon$ satisfies:
$N^{\prime}(\varepsilon)=O\left(\kappa(PA)\ln\left(\tfrac{1}{\varepsilon}\right)\right),$ (19)
Moreover, if $\mathcal{F}$ minimizes $\mathcal{L}_{\text{DATA}}$ this necessarily implies $\kappa(PA)=1\leq\kappa(A)$. Consequently, the number of steps is effectively reduced, i.e., $N^{\prime}(\varepsilon)\ll N(\varepsilon)$.
Proof.
Since $\mathcal{F}$ behaves like its linearization $P$, the gradient descent update becomes (refer to proof of Theorem 2 for steps):
$\Theta_{l+1}=\Theta_{l}-\eta P(A\Theta_{l}-b).$
Let $\Theta^{\ast}$ be the optimal parameter vector minimizing $\mathcal{L}_{\textnormal{PDE}}$. Then, the difference between the parameter vector at iteration $l$ and the optimal parameter vector is:
$\Theta_{l+1}-\Theta^{\ast}=\Theta_{l}-\Theta^{\ast}-\eta PA(\Theta_{l}-\Theta^{\ast})=(\text{Id}-\eta PA)(\Theta_{l}-\Theta^{\ast}).$
By recursively applying this update until the final step $L$, we obtain:
$\Theta_{L}-\Theta^{\ast}=(\text{Id}-\eta PA)^{L}(\Theta_{0}-\Theta^{\ast}).$
Since we have multiple data points, each with its own parameter vector, we consider the concatenation when necessary. Let’s introduce $\Xi_{l}$ as the matrix whose columns are the parameter vectors:
$\Xi_{L}=[\Theta_{L}^{(1)},\Theta_{L}^{(2)},\dots,\Theta_{L}^{(m)}].$
Similarly, $\Xi^{\ast}$ contains the optimal parameter vectors for each data point. The update for all data points can be written collectively:
$\Xi_{L}-\Xi^{\ast}=(\text{Id}-\eta PA)^{L}(\Xi_{0}-\Xi^{\ast}).$
Since $\mathcal{F}$ minimizes $\mathcal{L}_{\text{DATA}}$, we have $\Xi_{L}=\Xi^{\ast}$, implying:
$(\text{Id}-\eta PA)^{L}(\Xi_{0}-\Xi^{\ast})=0.$
Given that the values of $\Xi_{0}$ are iid and sampled randomly from a continuous distribution, because the set of singular matrices has measure zero, the square matrix $(\Xi_{0}-\Xi^{\ast})(\Xi_{0}-\Xi^{\ast})^{\top}$ is full rank (i.e., invertible), with probability 1. Thus, the only way for the above equality to hold is if:
$(\text{Id}-\eta PA)^{L}=0.$
This means $\text{Id}-\eta PA$ is nilpotent of index $L$. Consequently, all eigenvalues of $\text{Id}-\eta PA$ are zero, implying that all eigenvalues of $PA$ are equal to $\dfrac{1}{\eta}$, leading to $\kappa(PA)=\lambda_{\max}(PA)/\lambda_{\min}(PA)=1$, which is the optimal condition number. Referring to the convergence analysis in Theorem 2, we have:
$N^{\prime}(\varepsilon)\leq\frac{\kappa(PA)}{c}\ln\left(\frac{\|\Xi_{0}-\Xi^{\ast}\|_{2}}{\varepsilon}\right).$
Which directly implies
$N^{\prime}(\varepsilon)=O\left(\kappa(PA)\ln\left(\tfrac{1}{\varepsilon}\right)\right),$ (20)
With $\kappa(PA)=1$, this leads us to the desired result:
$N^{\prime}(\varepsilon)=O\left(\ln\left(\tfrac{1}{\varepsilon}\right)\right)\ll O\left(\kappa(A)\ln\left(\tfrac{1}{\varepsilon}\right)\right)=:N(\epsilon),$
Thus, the number of iterations required is significantly reduced compared to the case without the neural network preconditioner.
Discussion
The convergence proofs for our method fundamentally rely on the assumption of linearity in the underlying problems. It is important to note that the theoretical analysis does not extend to non-linear cases. Consequently, for non-linear scenarios, the theory should be viewed primarily as a tool for building intuition or providing motivation, rather than a definitive proof. This is due to the lack of established methods for rigorously studying the non-linear regime, as no known results currently address such cases. Under these conditions: